hi friends !!! can anyone explain the The Ideal Gas Equation Reply ASAP...

hi friends !!!


can anyone explain the The Ideal Gas Equation


Reply ASAP...

Grade:Upto college level

1 Answers

vijay narang
15 years ago

The relationships between volumes of gases under constant temperature and pressure were established by Boyle, Charles.

Boyles law was first published in 1662. It states

The pressure times the volume of a gas is equal to some constant k. PV= k.

Charles/Gay-Lussac Law, states that

At constant pressure, the volume of a given mass of an ideal gas increases or decreases by the same factor as its temperature (in kelvin) increases or decreases.

P1/T1 = P2/T2

Both of these laws can be combined to give an overall equation which describes the behaviour of gases.

P1V1/T1 = P2V2/T2

Avogadro's Law, states that the volume V is related to the number of moles of gas, n by

V/n = k

P1V1/T1n = P2V2/T2n = k
, where k is a constant.

Which is more familiar as

PV= nRT

The constant k has been replaced by R the molar gas constant.

Where P is the pressure, V is the volume, n is number of moles of gas present, R is the molar gas constant has a value very close to 8.31 m2 kg s-2 K-1 mol-1 and T is the temperature in Kelvin.